For normal incidennce at the air dielectric interface with dielectri constant Er=4. Find the fraction of energy reflected into the air?
1
Expert's answer
2020-05-05T18:44:12-0400
"R=\left(\frac{n_1-n_2}{n_1+n_2}\right)^2"

"n_1=1, n_2=\sqrt{E_r}=2"

"R=\left(\frac{1-2}{1+2}\right)^2=\frac{1}{9}"